The locus of the point from which two tangents to y^2=4ax make an angle A to each other is ??
THANKS IN ADVANCE..

Grade:

2 Answers

Vikas TU
11 years ago

You shud remmebr the formuale =

the angle substended  by the two tangents to a parabola y^2=4ax is

given by:

y^2 - 4ax = (x + a)2 . tan2A

Thus, it is only and can be the locus of the point from which two tangents to y^2=4ax make an angle A to each other.

 

 

Plz Approve!

:)

vivek kumar
10 years ago

apply ss1=t2 and then find angle between them and equate to a